OpenStudy (anonymous):

Multiply and Simplify: 3√ 6 × 2√ 12

OpenStudy (saifoo.khan):

\[36 \sqrt2\]

OpenStudy (anonymous):

\[3\sqrt{6}*2\sqrt{12} = 6\sqrt{72}=6\sqrt{36*2} = 6*6\sqrt{2} = 36 \sqrt{2}\]